OpenStudy (anonymous):

The mass of the ice cube is 17.6 g, and its initial temperature is -11.9 °C. The water resulting from the melted ice reaches the temperature of your skin, 29.9 °C. How much heat is absorbed by the ice cube and resulting water?

OpenStudy (aaronq):

you have to do this in a 3 step process, using equations: (1) \(q=H_{fusion}*m\) (2) \(q=m*C_p*\Delta T\)  Region A: heating up the ice from -11.9 to 0 degrees, use equation (2), and appropriate specific heat capacity. Region B: melting the ice (to water), use equation (1). Region C: heating up the water, use equation (2), use appropriate specific heat capacity.

OpenStudy (anonymous):

afterwards do you add the q values together, right?

OpenStudy (aaronq):

oh yes ! sorry, i forgot to mention that lol

OpenStudy (anonymous):

this is what i did Region A: (17.6)*(2.087)*(11.9)=437.101 Region B: I'm actually not sure but it says the enthalpy of fusion is 6010 j/mol and this is @ 1 mol Region C: (17.6)(4.184)*(29.9)=2199.68 and add the J Is this right?

OpenStudy (aaronq):

yeah thats right. for B since \(H_{fusion}\) is given in moles you have to convert the mass to moles. so (1) should be \(q=H_{fusion}*n\)

OpenStudy (anonymous):

okay cool thanks a lot! this helped

OpenStudy (aaronq):

no problem ! glad it did